what mass of Na2CO3 can be made if 1.52 g of NaHCO3 is reacted.

2NaHCO3(s)  →  Na2CO3(s) + H2O(l) + CO2(g)



1
Expert's answer
2020-11-09T13:52:18-0500

According to the equation of the reaction

2NaHCO3(s)  →  Na2CO3(s) + H2O(l) + CO2(g),

when 2 moles of NaHCO3 react, 1 mole of Na2CO3 is formed:

"\frac{n(NaHCO_3)}{2} = n(Na_2CO_3)" .

The number of the moles in 1.52 g of NaHCO3 (M = 84.01 g/mol) is :

"n(NaHCO_3) = \frac{m}{M} = \frac{1.52 \text{ g}}{84.01\text{ g/mol}} = 0.0181" mol.

Therefore, the number of the moles of Na2CO3 formed is:

"n(Na_2CO_3) = \frac{0.0181\text{ mol}}{2} = 0.0090" mol.

Finally, the mass of Na2CO3 produced (M = 105.99 g/mol) is:

"m(Na_2CO_3) = n·M = 0.0090\text{ mol}·105.99\text{ g/mol}"

"m(Na_2CO_3) =0.9588" , or 0.96 g.

Answer: 0.96 g of Na2CO3 can be made if 1.52 g of NaHCO3 reacted.